A block of mass m is connected to a spring of force constant k whose other is connected to a fixed wallInitially block is at rest and spring has natural length A constant force F is applied on the block away from the fixed wallThe max Speed of the block will be

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Let x be extension of spring at maximum speed, F=kx, x=F/k, max kinetic energy of block= Fx-(1/2)kx2

=F2/2k =mv2/2, thus maximum speed= v=F/√(mk)